# Aioli: A unified optimization framework for language model data mixing

### v1

@@ 6.2 Restricted Setting
Motivation.
We introduce the restricted setting because practitioners may not have the resources or desire to complete multiple full training runs to learn proportions, especially as recent LLMs are trained for longer and on more data [44]. As a result, practitioners may only use data mixing methods on shortened runs, producing learned proportions that may be suboptimal on the full run. We study if Aioli is able to improve performance by dynamically adjusting previously learned proportions throughout the full training run.
Setup.
We allow all existing methods only $0.5S$ additional training steps to learn the mixture proportions. This requires methods to learn $\bm{p}^{\text{method}}$ over shorter runs of $S_{\text{method}}$ steps each. For instance, grid search will conduct $10$ runs of length $S/20$ (see Table 11). We evaluate each method by using $\bm{p}^{\text{method}}$ learned from shorter runs to train the model on the full run of $S$ steps. We use Aioli to dynamically adjust each $\bm{p}^{\text{method}}$ throughout the full run. That is, for each existing method, we run Aioli with $\bm{p}^{\text{init}}=\bm{p}^{\text{method}}$ and $S_{\text{init}}=S_{\text{method}}$, referring to this as Aioli +method.
@@
DoReMi $-0.307$ $5.303$ $-0.217$ $-0.393$ $6.898$ $0.123$ 3 2
DoGE $0.419$ $0.184$ $-0.678$ $1.843$ $0.604$ $0.809$ 1 1
Aioli $-0.205$ $-0.340$ $-0.439$ $-0.096$ $-0.196$ $-0.240$ 6 0
Results.
In Table 3, we find that adding Aioli to any existing method that learns proportions over shorter runs improves average test perplexity per group in 28 out of 30 settings, by an average of 1.202 and a maximum of 12.012 points. Furthermore, Aioli can help methods that initially underperform stratified sampling surpass it, such as DoGE across all settings. In some settings, such as Books/StackExchange, Aioli improves methods that already outperform stratified sampling. This shows that Aioli can enhance a wide variety of mixture proportions, regardless of their initial performance. For the two cases where Aioli underperforms the base method, the base method already outperforms stratified, and adding Aioli maintains this trend, worsening perplexity by at most 0.025 points.
@@ 7 Related Work
Many data mixing methods utilize extra training runs to learn the static mixture proportions before the final training run. This allows us to view data mixing as a hyperparameter optimization problem in $\bm{p}$. [72] and [37] mitigate the inefficiency of grid search in higher dimensions by combining it with data mixing laws to impose additional structure. However, both grid search and these offline methods can have poor performance when $\bm{p}$ is searched for or fitted on shorter runs, as in the restricted setting. To understand these results, we note that many popular hyperparameter optimization methods carefully control truncation, and some runs are allowed to continue longer than others [35, 56, 18]. Thus, generic hyperparameter optimization methods may eventually prove effective for tuning data mixes.
@@ 8 Discussion
We propose an optimization framework, LMO, that can express several existing data mixing methods using their implicit mixing law. Using LMO, we show that linear dynamic mixing laws are well-specified in capturing the true loss-proportion relationship across datasets, but existing online methods still perform poorly on some datasets because their parameters are inaccurate. This insight inspires Aioli, whose performance gains are rooted in its ability to estimate parameters $A^{t}$ of the linear dynamic mixing law throughout training.
Limitations and Future Work.
Aioli introduces extra inference cost during training via the repeated evaluations in LearnParams (Alg. 2). This overhead can be reduced or completely mitigated by using the training loss as a proxy for the validation loss, computing $L_{\text{val}}$ over a subset of $D_{\text{val}}$, or by using each $A^{t}$ for longer (equivalently, decreasing $T$).
The LMO framework itself is an invitation for future work. We hope that LMO identifies clear axes of improvement for data mixing (the parameterization of the mixing law, parameter estimation, and how to solve for $\bm{p}$) and inspires the next generation of principled data mixing methods. Another direction for future work is understanding how to define data groups and how data group partitions impact data mixing. For example, C4 is a subset of CommonCrawl, so repartitioning these datasets into more disjoint groups could result in better data mixing performance. Furthermore, it is unclear if linear mixing laws results from some specific property of the data groups we studied. It would be interesting to consider in future work how often linear mixing laws hold, and how to exploit potential non-linearities in data mixing.
@@ 8.1 Reproducibility Statement
See Appendix B.2 for the full proofs on how to express Skill-it, DoReMi, and DoGE using the LMO framework. See Appendix C for details on how to reproduce our analyses of mixing law parametrization validity, $A^{t}$ parameter fit, and assessing whether greedy optimization is sufficient for data mixing. Finally, to reproduce the experimental results, please see Appendix D.
@@ 9 Acknowledgments
We thank Sabri Eyuboglu, Neel Guha, Ben Viggiano, Dan Biderman, Dan Fu, Michael Wornow, Jon Saad-Falcon, Alyssa Unell, Owen Dugan, Jerry Liu, and Gautam Machiraju for their feedback. We thank NYU HPC and Stanford NLP for providing compute and research support. This research project has benefited from the Microsoft Accelerate Foundation Models Research (AFMR) grant program.
@@
In Appendix A, we provide a glossary of notation used in the paper. In Appendix B, we discuss how additional data mixing methods are related to the LMO framework and provide proofs that existing methods can be expressed in our framework. In Appendix C, we provide additional results on our analysis of existing data mixing methods. In Appendix D we provide additional details for our results in Section 6, and in Appendix E we provide additional results, including downstream evaluation and ablations.
@@ B.1 Additional existing methods
RegMix [37] conducts many training runs on smaller models at shorter scales. Similar to DML [72], a regression model is fit to these runs and used to predict mixture proportions for a longer run on a larger model. They consider using a linear regression model, i.e., the mixing law $L_{\text{val},i}(\bm{p})=c_{i}-\sum_{j=1}^{m}A_{ij}p_{j}^{t}$, but find that the $R^{2}$ is relatively low (0.87). Instead, their main approach uses LightGBM, a tree-based gradient boosting approach, i.e., using an ensemble of non-linear decision trees as a mixing law. We note that Aioli could be used in conjunction with RegMix to facilitate better transfer across model size and training duration in their settings, an exciting direction for future work.
@@ B.2.1 Background on Exponentiated Gradient Descent
We provide background on exponentiated gradient descent (EGD) taken from Kakade [29]. In EGD, we have a sequence of decisions $w^{1},\dots,w^{T}$, where $w^{t}=[w_{1}^{t},\dots,w_{m}^{t}]\in\triangle^{m}$. We also have a sequence of cost functions $c^{1},\dots,c^{T}:\triangle^{m}\rightarrow\mathbb{R}$. To minimize the total cost $\sum_{t=1}^{T}c^{t}(w^{t})$, the EGD update rule sets $w^{0}=\text{Unif}(m)$, and updates according to $w_{j}^{t+1}=\frac{w_{j}^{t}\exp(-\eta\triangledown_{j}c^{t}(w^{t}))}{Z_{t}}$. $Z_{t}$ ensures that $w^{t+1}\in\triangle^{m}$, $\eta$ is a step size, and $\triangledown_{j}c^{t}(w^{t})$ denotes $\frac{\partial c^{t}(w^{t})}{\partial w_{j}^{t}}$. EGD is known to have certain regret guarantees on the value of costs incurred by playing $w^{1},\dots,w^{T}$ versus always playing the best fixed point in hindsight, i.e., the quantity $\sum_{t=1}^{T}c^{t}(w^{t})-\inf_{w\in\triangle^{m}}\sum_{t=1}^{T}c^{t}(w)$. We now are ready to prove Lemma 1.
@@
The cost function at each timestep in our setting is $\sum_{i=1}^{m}L_{\text{val},i}^{t+1}(\bm{p})$, and the decision we make is $p^{t}$. The mixing law constraint in (2) with $\sigma=\text{Id}$ is $L_{\text{val},i}^{t+1}(\bm{p})=c_{i}^{t}-b_{i}^{t}\sum_{j=1}^{m}A_{ij}^{t}p_{j}^{t}$ for all $i\in[m]$, so our objective function (1) can be written as
@@ B.2.2 Proof of Theorem 1
We note that [13] explicitly specifies their mixing law in equation 2 of their paper, along with the same objective function as ours in the LMO framework. ∎
@@
We comment on the fact that DoReMi’s proxy model is trained with a DRO (distributionally robust optimization) min-max objective, namely, $\text{minimize}_{f}\;\text{maximize}_{p}\sum_{i=1}^{m}p_{i}L_{\text{train},i}^{T+1}(f)$. This objective, which differs from our data mixing objective, yields $p^{t}$ gradient ascent and $f^{t}$ gradient descent updates. However, we are still able to express this training procedure in the LMO framework, since our claim is: if we assume that the $L_{\text{val},i}^{t+1}=L_{\text{val},i}^{t}(\bm{p})-b^{t}\sum_{j=1}^{m}A_{ij}^{t,\text{DRM}}p_{j}^{t}$ mixing law captures the relationship between $L_{\text{val}}^{t}$ and $p^{t}$, then training according to the DoReMi proxy run should not only guide $f$ and $\bm{p}$ to optimize the DRO objective, but also to optimize the average validation loss per group.
@@
Let each group $i$ consist of samples $x$ from the distribution $\mathcal{D}_{i}$, where $D_{\text{train}}\sim\mathcal{D}_{i}$ and let $\tilde{L}_{\text{train, i}}(f)=\mathbb{E}_{x\sim\mathcal{D}_{i}}\left[\ell(f,x)\right]$ be the population-level loss on group $i$, where $\ell(f,x)$ is $f$’s loss on sample $x$.
If a batch is uniformly sampled, each group has $B/m$ samples. We can then write $L_{B,i}^{t}(f,\text{Unif}(m))=\sum_{k=1}^{B/m}\ell(f,x_{k}^{i})$, where $x_{k}^{i}$ is the $k$th sample of group $i$ in the batch. Then,
@@
Appendix C Analysis Details and Additional Results
@@ C.1.1 Additional parameterization experiments
Parameterization across checkpoints. We investigate whether the log-linear static and linear dynamic mixing laws remain well-specified in later stages of training and on other datasets. To do so, we take various Pythia 160M checkpoints  [8], sweep mixing proportions, and fit the linear dynamic and log-linear static mixing laws. We train for $2000$ steps according to the learning rates and learning rate scheduler reported in [8]. We fit the static mixing law on full runs of $2000$ steps, and the linear dynamic mixing law at $t=500$, after which we do a training sweep over the next $500$ steps. In Tables 6 and 7, we find that the strong fit for log-linear static mixing laws continues to hold during pre-training at checkpoint 72K (roughly halfway through training Pythia-160M) and after pre-training, with an average $R^{2}$ of 0.982 and 0.991, respectively. However, the linear dynamic mixing law’s $R^{2}$ coefficient is lower, averaging 0.815 at checkpoint 72K and 0.830 at the end of pre-training. It may be interesting to further study if the dynamics of the loss-proportion relationship evolve in a structured way throughout training, or if these results are due to more noise in how models learn at later stages of training.
Parameterization across other sets of data groups. In Figure 4, we identify an example set of data groups that exhibits a non-linear relationship between loss and proportion: Books/C4 from SlimPajama. For these two data groups, we see that as the proportion of Books increases while C4 decreases, the loss on Books starts increasing past a certain $\bm{p}$, suggesting quite counterintuitively that performance on Books is optimized by allocating some proportion to C4. In this case, neither log-linear static or linear dynamic mixing laws have good fit to the proportion-loss relationship, as neither can represent the non-linearity. In particular, the average MSE and $R^{2}$ for the log-linear static mixing law is $0.003$ and $0.558$, respectively, and the average MSE and $R^{2}$ for the linear dynamic mixing law is $0.0002$ and $0.721$.
Fortunately, because these nonlinearities exist on the boundary of the simplex and tend to incur high loss, they tend to have little impact on the optimization of $\bm{p}$, which strives to minimize the average loss for the standard data mixing problem. For instance, we found that the optimal proportion according to Ye et al. [72]’s log-linear static mixing law on one random seed was $[0.176,0.824]$, and the true optimal from grid search was $[0.2,0.8]$. However, it is important to further investigate this non-linear phenomenon on additional data groups and training regimes, which we defer to future work.
@@
Figure 4: Top: Log-linear static mixing law fit on Books/C4 across 5 random seeds (validation loss). Bottom: Linear dynamic mixing law fit on Books/C4 on 1 random seed. Each color is a different initial mixture $p^{0}\in\mathcal{P}$ trained for $2000$ steps, and the fitting sweeps are done over $100$ additional steps.

### v2

@@ 6.2 Restricted Setting
Motivation. We introduce the restricted setting because practitioners may not have the resources or desire to complete multiple full training runs, especially as recent LLMs are trained for longer and on more data [45]. As a result, practitioners may only use data mixing methods on shortened runs, producing learned proportions that may be suboptimal on the full run. We study if Aioli is able to improve performance by dynamically adjusting previously learned proportions throughout the full training run.
Setup. We allow all existing methods up to $0.5S$ additional training steps to learn the mixture proportions. This requires methods to learn $\bm{p}^{\text{method}}$ over shorter runs of $S_{\text{method}}$ steps each. For instance, grid search will conduct $10$ runs of length $S/20$ (see Table 11). We evaluate each method by using $\bm{p}^{\text{method}}$ learned from shorter runs to train the model on the full run of $S$ steps. We use Aioli to dynamically adjust each $\bm{p}^{\text{method}}$ throughout the full run. That is, for each existing method, we run Aioli with $\bm{p}^{\text{init}}=\bm{p}^{\text{method}}$ and $S_{\text{init}}=S_{\text{method}}$, referring to this as Aioli +method.
@@
DoReMi $-0.307$ $5.303$ $-0.217$ $-0.393$ $6.898$ $0.703$ 3 2
DoGE $0.419$ $0.184$ $-0.678$ $1.843$ $0.604$ $0.949$ 1 1
Aioli $-0.205$ $-0.340$ $-0.439$ $-0.226$ $-0.196$ $-0.240$ 6 0
Results. In Table 3, we find that adding Aioli to any existing method that learns proportions over shorter runs improves average test perplexity per group in 28 out of 30 settings, by an average of 1.202 and a maximum of 12.012 points. Furthermore, Aioli can help methods that initially underperform stratified sampling surpass it, such as DoGE across all settings. In some settings, such as Books/StackExchange, Aioli improves methods that already outperform stratified sampling. This shows that Aioli can enhance a wide variety of static proportions, regardless of their initial performance. For the two settings where Aioli underperforms the base method, the base method already outperforms stratified, and adding Aioli maintains this trend, worsening perplexity by at most 0.025 points.
@@ 7 Related Work
Many data mixing methods utilize extra training runs to learn the static mixture proportions before the final training run. This allows us to view data mixing as a hyperparameter optimization problem in $\bm{p}$. [72] and [37] mitigate the inefficiency of grid search in higher dimensions by combining it with data mixing laws to impose additional structure. However, both grid search and these offline methods can have poor performance when $\bm{p}$ is searched for or fitted on shorter runs, as in the restricted setting. To understand these results, we note that many popular hyperparameter optimization methods carefully control truncation, and some runs are allowed to continue longer than others [35, 57, 19]. Thus, generic hyperparameter optimization methods may also prove effective for tuning data mixes.
@@ 8 Discussion
We introduce the LMO framework, which unifies existing data mixing methods by viewing them as solutions to a common optimization problem involving an implicit method-dependent mixing law. Using this framework, we find that existing methods perform poorly on some datasets due to inaccurate mixing law parameters. This insight inspires Aioli, whose performance gains are rooted in its ability to estimate parameters $A^{t}$ of the linear dynamic mixing law throughout training.
Limitations and Future Work Aioli incurs extra inference cost via the repeated evaluations in LearnParams (Alg. 2). This can be reduced by computing $L_{\text{val}}$ over a subset of $D_{\text{val}}$, and by using each $A^{t}$ for longer (decreasing $T$). Another direction is understanding the role of data group partitions. For example, C4 is a subset of CommonCrawl, and it is unclear if disjoint groups could improve performance.
The LMO framework itself is an invitation for future work. It shows that data mixing methods can be improved and analyzed by studying their assumptions on how models learn from data. By exposing such assumptions, LMO identifies key axes for improvement (mixing law parameterization, parameter estimation, and how to solve for $\bm{p}$), which we hope will inspire new principled data mixing methods.
@@ 8.1 Reproducibility Statement
See Appendix B.2 for the full proofs on how to express Skill-it, DoReMi, and DoGE using the LMO framework. See Appendix C for details on how to reproduce our analyses of mixing law parametrization validity, $A^{t}$ parameter fit, and assessing whether greedy optimization is sufficient for data mixing. Finally, to reproduce the experimental results, please see Appendix E.
@@ 9 Acknowledgments
We thank Sabri Eyuboglu, Neel Guha, Ben Viggiano, Dan Biderman, Dan Fu, Michael Wornow, Jon Saad-Falcon, Alyssa Unell, Owen Dugan, Jerry Liu, and Gautam Machiraju for their feedback. We thank Stanford NLP for providing compute and research support. This work was supported in part through the NYU IT High Performance Computing resources, services, and staff expertise. This research project has benefited from the Microsoft Accelerate Foundation Models Research (AFMR) grant program.
@@
In Appendix A, we provide a glossary of notation used in the paper. In Appendix B, we discuss how additional data mixing methods are related to the LMO framework and provide proofs that existing methods can be expressed in our framework. In Appendix C, we provide additional results on our analysis of existing data mixing methods. In Appendix E we provide additional details for our results in Section 6, and in Appendix F we provide additional results, including downstream evaluation and ablations.
@@ B.1 Additional existing methods
RegMix [37] conducts many training runs on smaller models at shorter scales. Similar to DML [72], a regression model is fit to these runs and used to predict mixture proportions for a longer run on a larger model. They consider using a linear regression model, i.e., the mixing law $L_{\text{val},i}(\bm{p})=c_{i}-\sum_{j=1}^{m}A_{ij}p_{j}^{t}$, but find that the $R^{2}$ is relatively low (0.87). Instead, their main approach uses LightGBM, a tree-based gradient boosting approach, i.e., using an ensemble of non-linear decision trees as a mixing law. We note that Aioli could be used in conjunction with RegMix in their settings, an exciting direction for future work.
@@ B.2.1 Background on Exponentiated Gradient Descent
We provide background on exponentiated gradient descent (EGD) taken from Kakade [29]. In EGD, we have a sequence of decisions $w^{1},\dots,w^{T}$, where $w^{t}=[w_{1}^{t},\dots,w_{m}^{t}]\in\triangle^{m}$. We also have a sequence of cost functions $c^{1},\dots,c^{T}:\triangle^{m}\rightarrow\mathbb{R}$. To minimize the total cost $\sum_{t=1}^{T}c^{t}(w^{t})$, the EGD update rule sets $w^{0}=\text{Unif}(m)$, and updates according to $w_{j}^{t+1}=\frac{w_{j}^{t}\exp(-\eta\triangledown_{j}c^{t}(w^{t}))}{Z_{t}}$. $Z_{t}$ ensures that $w^{t+1}\in\triangle^{m}$, $\eta$ is a step size, and $\triangledown_{j}c^{t}(w^{t})$ denotes $\frac{\partial c^{t}(w^{t})}{\partial w_{j}^{t}}$. EGD is known to have certain regret guarantees on the value of costs incurred by playing $w^{1},\dots,w^{T}$ versus always playing the best fixed point in hindsight: $\sum_{t=1}^{T}c^{t}(w^{t})-\inf_{w\in\triangle^{m}}\sum_{t=1}^{T}c^{t}(w)$.
We now are ready to prove Lemma 1.
@@
The cost function at each timestep in our setting is $\sum_{i=1}^{m}L_{\text{val},i}^{t+1}(\bm{p})$, and the decision we make is $p^{t}$. The mixing law constraint in (2) with $\sigma=\text{Id}$ is $L_{\text{val},i}^{t+1}(\bm{p})=c_{i}^{t}-b_{i}^{t}\sum_{j=1}^{m}A_{ij}^{t}p_{j}^{t}$ for all $i\in[m]$, so our objective (1) can be written as
@@ B.2.2 Proof of Theorem 1
We note that [14] explicitly specify their mixing law in equation 2 of their paper, along with the same objective function as ours in the LMO framework. ∎
@@
We comment on the fact that DoReMi’s proxy model is trained with a DRO (distributionally robust optimization) min-max objective, namely, $\text{minimize}_{f}\;\text{maximize}_{p}\sum_{i=1}^{m}p_{i}L_{\text{train},i}^{T+1}(f)$. This objective, which differs from our data mixing objective, yields the $p^{t}$ gradient ascent and $f^{t}$ gradient descent updates. However, we are still able to express this training procedure in the LMO framework, since our claim is: if we assume that the $L_{\text{val},i}^{t+1}=L_{\text{val},i}^{t}(\bm{p})-b^{t}\sum_{j=1}^{m}A_{ij}^{t,\text{DRM}}p_{j}^{t}$ mixing law captures the relationship between $L_{\text{val}}^{t}$ and $p^{t}$, then training according to the DoReMi proxy run should not only guide $f$ and $\bm{p}$ to optimize the DRO objective, but also to optimize the average validation loss per group.
@@
Let each group $i$ consist of samples $x$ from the distribution $\mathcal{P}_{i}$, and let $\tilde{L}_{\text{train, i}}(f)=\mathbb{E}_{x\sim\mathcal{P}_{i}}\left[\ell(f,x)\right]$ be the population-level loss on group $i$, where $\ell(f,x)$ is $f$’s loss on sample $x$.
If a batch is uniformly sampled, each group has $B/m$ samples. We can then write $L_{B,i}^{t}(f,\text{Unif}(m))=\sum_{k=1}^{B/m}\ell(f,x_{k}^{i})$, where $x_{k}^{i}$ is the $k$th sample of group $i$. Then,
@@
Appendix C Analysis Details
@@ C.1.1 Additional parameterization experiments
Parameterization across checkpoints.
We investigate whether the log-linear static and linear dynamic mixing laws remain well-specified in later stages of training and on other datasets. To do so, we take various Pythia 160M checkpoints  [8], sweep mixing proportions, and fit the linear dynamic and log-linear static mixing laws. We train for $2000$ steps according to the learning rates and learning rate scheduler reported in [8]. We fit the static mixing law on full runs of $2000$ steps, and the linear dynamic mixing law at $t=500$, after which we do a training sweep over the next $500$ steps. In Tables 6 and 7, we find that the strong fit for log-linear static mixing laws continues to hold during pre-training at checkpoint 72K (roughly halfway through training Pythia-160M) and after pre-training, with an average $R^{2}$ of 0.982 and 0.991, respectively. However, the linear dynamic mixing law’s $R^{2}$ coefficient is lower, averaging 0.815 at checkpoint 72K and 0.830 at the end of pre-training. It thus may be interesting to further study if the dynamics of the loss-proportion relationship evolve in a structured way throughout training, or if these results are due to more noise in how models learn at later stages of training.
Parameterization across other sets of data groups.
In Figure 4, we identify an example set of data groups that exhibits a non-linear relationship between loss and proportion: Books/C4 from SlimPajama. For these two data groups, we see that as the proportion of Books increases while C4 decreases, the loss on Books starts increasing past a certain $\bm{p}$, suggesting quite counterintuitively that performance on Books is optimized by allocating some proportion to C4. In this case, neither log-linear static or linear dynamic mixing laws have good fit to the proportion-loss relationship, as neither can represent the non-linearity. In particular, the average MSE and $R^{2}$ for the log-linear static mixing law is $0.003$ and $0.558$, respectively, and the average MSE and $R^{2}$ for the linear dynamic mixing law is $0.0002$ and $0.721$.
Fortunately, because these nonlinearities exist on the boundary of the simplex and tend to incur high loss, they tend to have little impact on the optimization of $\bm{p}$, which strives to minimize the average loss. For instance, we found that the optimal proportion according to Ye et al. [72]’s log-linear static mixing law on one random seed was $[0.176,0.824]$, and the true optimal from grid search was $[0.2,0.8]$. However, it is important to further investigate this non-linear phenomenon on additional data groups and training regimes, which we defer to future work.
@@
Figure 4: Top: Log-linear static mixing law fit on Books/C4 across 5 random seeds. Bottom: Linear dynamic mixing law fit on Books/C4 on 1 random seed. Each color is a different initial mixture $p^{0}\in\mathcal{P}$ trained for $2000$ steps, and the fitting sweeps are done over $100$ additional steps.
@@ C.1.2 Parameterization on instruction-tuning mixtures
Checking for interactions among groups.
It is natural to ask whether a linear mixing law is sufficient to model how mixing proportions affect the loss. In linear regression, such assumptions are often evaluated using visual diagnostics called residual plots [43]. Residual plots graph the prediction error from each data point (the residuals) in order to reveal different kinds of structure. For example, it is common to plot the residual against the predicted value to check for nonlinearity.
Figure 5 shows several such residual plots for the dynamic mixing law experiments with 3 domains (Arxiv, Books, and Stackexchange). The figure checks for interactions when predicting Arxiv’s loss. The corresponding plots for the other domains look similar.
The top row visualizes the residuals inside the simplex. If strong interactions were present, then they would cause clustered patterns in the residuals—regions where the linear model consistently gives predictions that are too low or too high. Strong patterns do not seem apparent.
The bottom three rows plot the residuals against different interaction terms. A consistent trend in the residuals above or below zero would suggest the term captures a meaningful interaction. The scatter plots show no consistent trend. The first three charts on the bottom row hint that a small interaction could be present in those cases; however, it is difficult to say without larger samples. Considering the linear model’s excellent fit and high $R^{2}$, if such an interaction is present then it is likely small.
To summarize: the linear model seems sufficient. While we can not rule out the possibility of small interactions, the diagnostics do not reveal any major departures from linearity that might compel us to use a more complex model.
Figure 5: Residuals plots to check for interactions in the dynamic mixing law experiments with 3 domains (Arxiv, Books, and StackExchange). The target loss is Arxiv. Columns correspond to different initial mixing proportions. Data points show the (externally studentized) residuals of different mixing proportions after fitting the linear mixing law. Top row: Each point in the simplex corresponds to a different mixture of the 3 domains, with its color giving the residual’s value at that point (red is positive, blue is negative). Bottom 3 rows: each row shows the residual plotted against a different interaction term: $P_{1}P_{2}$, $P_{1}P_{3}$, and $P_{2}P_{3}$. Dotted gray lines show upper and lower 99% confidence limits for the residuals, assuming the linear regression assumptions hold.